OpenStudy (anonymous):

r=30000p-2000p^2 how do calculate the maximum revenue using calculus ??

hero (hero):

Use x = -b/2a

hero (hero):

b = 30000 a = -2000

hero (hero):

15000

hero (hero):

But we have to find y

OpenStudy (anonymous):

\[ r \prime = 30000 -4000 p\]

hero (hero):

I mean, r, not y, lol

OpenStudy (anonymous):

30000-4000 p =0 ===>p = 7.5 ===r(7.5) = 112500

hero (hero):

sorry p = 7.5

hero (hero):

r = 112,500 as raheen stated

OpenStudy (anonymous):

yes but how do you guys get 112500 i can only get it to equal 0

hero (hero):

There's two ways to find it...Either use p = -b/2a or take the derivative, set it equal to zero and solve for p. I recommend using p = -b/2a. It's easier

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

since the calculus metheod was required, one must solve by derivative method